Invoice Report

January – April 1994

NMPC Project 07-9199

High Performance Computing and Parallel Processing in the Utility Industry

Geoffrey Fox, Paul Coddington, Alok Choudhary, Salim Hariri, Michael Nilan, Sanjay Ranka

Northeast Parallel Architectures Center, Syracuse University

Work completed on this project for the period January 1 to April 30 1994:

• Improvements and further development of a parallel sparse matrix program for power systems problems.

• Work on development of a prototype parallel version of the EPRI ETMSP transient stability analysis program, in particular an interface to the parallel sparse matrix program.

• Development of a prototype version of an expert system to analyze transient stability output.

• Investigation into methods for porting the parallel programs to the IBM SP-1 parallel computer, and work to implement this port.

See attached financial summary for the expenditures for this period for the work described in this invoice, and for expenditures to-date for this project.

The work described here involved 100% of effort for 1 researcher and 1 software engineer, 20% of effort for 1 faculty member, 6% of effort for 1 faculty members, 5% of effort for 2 additional faculty members, and 65% of effort for 5 graduate students.

Total including fringe benefits and overhead is $72,105.24 for this invoice, $200,000 to-date.